Suppose the average total cost of producing 9 sweaters is $65 and the marginal cost of producing the 10th sweater is $67. Everything else held constant, the average total cost of producing 10 sweaters will be _____ the average total cost of producing 9 sweaters.

a. greater than

b. the same as

c. less than
Answer #1

option a

average cost of producing 9 sweaters = $65

total cost of 9 sweaters = 65*9 = $585

marginal cost of 10th sweater = $67

total cost of 10 sweater s= $585 + $67 = $652

average cost of 10 sweater s= 652/10 = $65.2

so average cost of producing 10 sweaters , $65.20 > average cost of 9 sweaters